If c ( q ) = 100 + 4 q 2 and q ∗ = 16, then we know that the market price must be:
Answer #1

Ans) market price must be: $ 128

At Profit maximization, P = MC

MC = dC/dQ = 8q

Given Q = 16 implies MC = 16*8 = 128 so price = 128